# A Differentiable Model of Nucleic Acid Dynamics

### v1

@@ Abstract
Developing physics-based models for molecular simulation requires fitting many unknown parameters to experimental datasets of diverse properties. Fitting such diverse data is typically carried out piecemeal, resulting in a “zoo” of disconnected but highly related models. Here, we leverage recently developed methods for computing low-variance gradient estimates with automatic differentiation to facilitate the extensible, reproducible, and transparent fitting of the physics-based DNA model oxDNA. We demonstrate simultaneous fitting to structural, mechanical, and thermodynamic data using a variety of techniques, including enhanced sampling, external forcing, cluster-move Monte Carlo schemes, and micro- and millisecond timescales, and fit oxDNA to a range of target properties, such as twist-stretch couplings and duplex melting temperatures. We also demonstrate how gradients can serve as an efficient means of sensitivity analysis through which we uncover the surprising importance of the cross-stacking interaction in oxDNA. Though grounded in oxDNA, this work provides a framework for extensible, community-driven force field development more generally that will yield rapidly improved models and subsequent physical insights for systems ranging from RNA to protein to their interactions with inorganic materials.
@@ 1 Introduction
Physics-based molecular simulations using coarse-grained potentials are a cornerstone of in silica biophysical modeling. When sufficiently accurate, they provide a unique way to study a biophysical system beyond available experimental data [1]. Such models have provided deep insights ranging from viral assembly [2, 3, 4, 5] to protein folding [6, 7, 8] to chromosomal organization [9, 10]. Recent successes in experimental nucleic acid nanotechnology [11, 12, 13] combined with the comparative simplicity of DNA and RNA relative to proteins renders DNA/RNA modeling an attractive and active area of research.
@@
Yet, while recent years have seen extraordinary progress in the accuracy of large scale machine learning (ML) models, from large language models [14, 15] to protein structure prediction [16], the accuracy of coarse-grained potentials has stagnated. The machine learning examples clearly demonstrate that increased model accuracy unlocks enormous value. Recently, Donoho hypothesized [17] that frictionless reproducibility, the ease of reproducing and iteratively improving models, is the major factor leading to dramatic ML model improvement. Within ML, model parameters are almost universally fit via gradient-based optimization, datasets are clearly partitioned into train, validation, and test subsets, objective functions are clearly-defined, and trained models and fitting procedures are often publicly available via popular open-source tools such as GitHub and Hugging Face. These standardized and transparent fitting procedures enable rapid technological innovation.
Here, we demonstrate how this paradigm can be applied to fitting coarse-grained, physics-based potentials. We apply differentiable programming, the backbone of machine learning algorithms whereby gradients are computed via automatic differentiation (AD), to an exemplar of coarse-grained molecular modeling: nucleic acid force field development.
Despite resounding successes for particular coarse-grained DNA and RNA models [18, 19, 20, 21, 22, 23], both types of force field development are plagued by suboptimal parameter fitting methods that are often opaque and hard to reproduce. Force field parameters are typically the result of heroic fitting procedures with significant hand-design [24, 18, 25, 26, 27, 28, 29, 30]. Even force fields that have been fit through comparatively more principled approaches such as Iterative Boltzmann Inversion [31, 24], maximum likelihood methods [32], or Newton Inversion [33] often use piecemeal approaches to parameter optimization: each type of interaction (e.g. hydrogen bonding, excluded volume, stacking, etc) is treated individually – often with differing parameterization schemes – without simultaneous optimization of all parameter types. As such, existing models cannot be easily adapted to model new effects (e.g. non-canonical base-pairing, electrostatics) or new experimental data. Often, entirely new models are invented to treat experimental inconsistencies without knowing whether this is truly required given the initial model parameterization.
A case in point is the recent MADna model of Assenza and Pérez [24], developed to model mechanical properties that are poorly captured by existing models (e.g. oxDNA, 3SPN). Rather than fine-tuning an existing, otherwise successful DNA model, Assenza and Pérez were required to develop an entirely new model. Indeed, expanded parameterizations of DNA/RNA models often appear years after the original model release [34, 35, 36, 37, 38], attesting to the difficulty and laboriousness of refining models where parameterization is nontransparent and unsystematic. The ultimate result is a fragmented “zoo” of models that each require months or years of trial-and-error parameter fitting despite attempting to represent similar datasets.
Figure 1: The formulation of force field parameter fitting as an optimization problem akin to neural network training. The choice of functional form is analogous to fixing a neural network architecture, where geometric parameters control the relative positions of beads in a rigid body and Hamiltonian parameters describe the interactions between these beads. The objective function is formulated as a function of ground truth data, typically derived from experiment, and is evaluated by means of forward simulations. The oxDNA model was originally fit to three classes of experimental data: structural (e.g. helical pitch and diameter), mechanical (e.g. persistence length $L_{ps}$), and thermodynamic (e.g. melting temperatures $T_{m}$). The depiction of the oxDNA architecture is adapted from Ref. [39].
Here, we focus on solving these issues within oxDNA, a coarse-grained model with a physics-based and interpretable potential energy function. The model has enjoyed extensive validation against experiment [40] and is therefore widely employed in the DNA nanotechnology community for both (i) deriving in silica biophysical insights in systems such as nanotweezers and gels [41, 42, 43], and (ii) complementing experimental studies in subjects ranging from strand displacement [44, 45] to DNA origami [46, 47]. By focusing on an experimentally-relevant model with a relatively complex and expressive energy function, we are forced to confront the three main issues of traditional differentiable MD. First, computing the gradient through long unrolled simulations imposes a significant memory overhead as well as numerical instabilities [48, 49]. Second, many fundamental algorithms in molecular simulation are not end-to-end differentiable (e.g. cluster-move Monte Carlo schemes). Third, decades of research have been dedicated to optimizing MD algorithms to increase the accessible timescales, but many of these optimizations are lost upon reimplementation in an automatic differentiation framework.
@@
We surmount these technical challenges using a novel method for stochastic gradient estimation [50, 51], making it possible to use gradient-descent to fit force field parameters to the three classes of experimental data to which oxDNA was originally fit: structural, mechanical, and thermodynamic. Each class of data poses a unique set of challenges that are not present in toy models previously studied with differentiable MD. For example, calculation of mechanical properties, such as persistence length or stretch modulus, typically requires micro- and millisecond-scale simulations to sample many uncorrelated states, line-fitting procedures to infer summary statistics, and the application of external forces and torques. Similarly, melting temperature calculations require particularly complex simulations, involving enhanced sampling, cluster-move Monte Carlo algorithms, and temperature extrapolation. Our approach allows us to combine the methods developed for each individual class of data to fit parameters to single objective functions that jointly target structural, mechanical, and thermodynamic properties, allowing for the first time a systematic, simultaneous parameter optimization for different loss functions and for different interaction terms in the potential energy function. We also demonstrate how the calculated gradients can be adapted for sensitivity analyses, enabling the rapid derivation of physical insights. Our results not only agree with existing literature, but also reveal new insights. As an example, we find cross-stacking interaction parameters are much more influential than naively expected.
We hope that this work serves as a launching point for what we believe will be paradigm shift in extensible, collaborative, community-driven force field development. We make all code for the parameter optimizations presented in this work available at the following GitHub repository: https://github.com/rkruegs123/jax-dna.
Figure 2: Traditional differentiable molecular dynamics does not scale to oxDNA, a representative experimentally-relevant force field. A. The traditional methods for stochastic gradient estimation in which gradients are directly calculated via a differentiable simulator (i.e. the pathwise estimator) or the gradients of (log) probabilities of individual trajectories are scaled by a reward defined for each trajectory (i.e. the score function estimator). Conventional differentiable molecular dynamics relies on the pathwise estimator (see Supporting Information for an evaluation of the score function estimator). B. The memory cost per pathwise estimator gradient calculation as a function of simulation length for a 60 bp duplex, the prototypical system for evaluating structural properties, revealing the maximum simulation length ($\sim$ 2500 steps) feasible on currently available hardware. C. The (log) average absolute value of the gradient of the RMSE of the measured pitch vs. the target pitch as a function of simulation length for simulations of a 60 bp duplex. Gradient rematerialization (also called gradient checkpointing) was used to permit gradient calculations at longer simulation lengths at the cost of runtime (see Supporting Information). D. A comparison of the time to simulate an eight base pair duplex for 2000 timesteps with and without gradient calculation.
@@ 2.1 Differentiable Force Field Implementation
To carry out gradient-descent on the parameters of oxDNA, we first implemented the oxDNA energy function in JAX, a state-of-the-art AD library for scientific computing [52]. There are several versions of oxDNA as additional features such as sequence-specificity and major-minor grooving have been added since the parameterization of the original model. While we implement all versions of oxDNA in our framework, here (unless otherwise stated) we focus our results on the sequence-averaged version of oxDNA2, the latest version of oxDNA. To simulate systems of DNA with our implementation, we use JAX-MD, an end-to-end differentiable molecular dynamics engine written in JAX [53]. We benchmark our implementation against the standalone oxDNA code [54, 55], involving both highly-optimized C++ and CUDA implementations, and find that our implementation is 2-3x slower (Figure S1).
Despite this performance cost, the JAX implementation offers two key advantages. First, since we write our energy function in an AD framework, forces are computed automatically from energy functions. This significantly reduces the barrier to experimenting with modified functional forms of the energy function, and also permits the application of arbitrary (continuous and differentiable) external forces and torques. Second, since the code can be compiled to a range of hardware accelerators, we can immediately run simulations on GPU and TPU without reimplementing the force field in a separate framework. Therefore, our JAX implementation can be seen as complementary to the existing standalone and LAMMPS [56] implementations of oxDNA: beyond permitting gradient-based optimization, it enables flexibility and extensibility for force field development.
We now in turn consider each of the three classes (i.e. structural, mechanical, and thermodynamic) of data used to fit force field parameterizations in oxDNA.
@@ 2.2 Structural Optimization
Structural properties are the most amenable to optimization via differentiable MD, since the simulations required to compute such properties are shortest in time and smallest in system size. As a first step, we fit a subset of oxDNA parameters to target pitch and propeller twist values. We sought to evaluate whether the pathwise estimator (Figure 2A), whereby gradients are computed directly through an unrolled numerical simualation, is sufficient for optimizing oxDNA parameters for structural properties. We find that the pathwise estimator has numerous numerical and computational limitations; the backwards calculation imposes significant time and memory overhead, and yields gradients whose magnitude increases sharply with simulation time (Figure 2B-D). The gradients calculated through short simulations for computing structural properties were too numerically unstable for optimization (Figure S2A). At the same time, we also evaluated the score function estimator for structural optimization, in which batches of simulations are performed and the probability of such simulations are increased or decreased according to the calculated observable (Figure 2A). This method also did not yield sufficient signal for optimization (Figure S2B).
Figure 3: Optimization of oxDNA parameters for target structural properties via trajectory reweighting. A. An overview of differentiable Monte Carlo (or “DiffTRE”). B. The average absolute value of the gradient as a function of simulation steps when using DiffTRE for stochastic gradient estimation. C. The pitch as a function of iteration when using the gradients calculated via DiffTRE for gradient-based optimization. D. The ten most significant parameters ranked by the absolute value of their gradients.
@@
As such, we turned to a recently developed method of stochastic gradient estimation that directly operates at the level of unnormalized steady state probability distributions (Figure 3A). This method was first introduced as a general method for differentiable Monte Carlo in the context of variational quantum Monte Carlo by Zhang et al. [50] and machine-learned potentials for molecular simulation [51], where it was termed DiffTRE, serving as a low-variance score function estimator that effectively computes the objective function as an integral over unnormalized probability distributions; see Materials and Methods for details. This estimator circumvents the memory and numerical issues associated with traditional differentiable MD by only differentiating the energy of reference states sampled from a set of simulations.
@@
We find that for fitting structural properties to oxDNA, DiffTRE provides a numerically stable, scalable, and efficient method for optimization via gradient descent (Figure 3B). Unlike the pathwise and score function estimators, this method yielded oxDNA parameters for specified pitch values within 50 iterations (Figure 3C), and the procedure was stable across various target pitch values. It is important to note that DiffTRE does not require that the sampling procedure itself is differentiable. This means we can optimize oxDNA parameters while using the standalone oxDNA implementation, a highly optimized implementation of the force field, rather than JAX-MD.
@@
For a physics-based model, the gradient estimates for a fixed parameter set contain information about the most important physical interactions for a target property, accessible by ranking parameters based on the magnitude of their gradient (Figure 3D). For pitch, we find that the most significant parameter is the equilibrium distance of the backbone sites, as previously found [37]. More surprisingly, we find that cross-stacking interactions are quite important: this interaction is most often implicated for controlling duplex melting temperatures given its stabilizing role [57, 58, 59]. Sensitivity analysis for propeller twist (Figure S3) also reveals the two most important parameters are involved in the cross-stacking interaction.
Finally, we turn to a sequence-structure pair deposited in Nanobase [60], optimizing the oxDNA parameters to better reproduce the gated DNA nanopore introduced in Ref. [61]. We can choose parameters to minimize the root mean square deviation (RMSD) between the simulated structure and the target structure (Figure S4).

### v2

# Fitting Coarse-Grained Models to Macroscopic Experimental Data via Automatic Differentiation

@@ Abstract
Developing physics-based models for molecular simulation requires fitting many unknown parameters to diverse experimental datasets. Traditionally, this process is piecemeal and difficult to reproduce, leading to a fragmented landscape of models. Here, we establish a systematic, extensible framework for fitting coarse-grained molecular models to macroscopic experimental data by leveraging recently developed methods for computing low-variance gradient estimates with automatic differentiation. Using a widely validated DNA force field as an exemplar, we develop methods for optimizing structural, mechanical, and thermodynamic properties across a range of simulation techniques, including enhanced sampling and external forcing, spanning micro- and millisecond timescales. We highlight how gradients enable efficient sensitivity analyses that yield physical insight. We then demonstrate the broad applicability of these techniques by optimizing diverse biomolecular systems, including RNA and DNA-protein hybrid models. We show how conflict-free gradient methods from multi-task learning can be adapted to impose multiple constraints simultaneously without compromising accuracy. This approach provides a foundation for transparent, reproducible, community-driven force field development, accelerating progress in molecular modeling.
@@ 1 Introduction
Physics-based molecular simulations using coarse-grained models are a cornerstone of in silica biophysical modeling. When sufficiently accurate, they provide a unique way to study a biophysical system beyond available experimental data [1]. Such models have provided deep insights ranging from viral assembly [2, 3, 4, 5] to protein folding [6, 7, 8] to chromosomal organization [9, 10].
@@
Recent years have seen extraordinary progress in the accuracy of large-scale machine learning (ML) models, from large language models [11, 12] to protein structure prediction [13]. Donoho hypothesized [14] that frictionless reproducibility, the ease of reproducing and iteratively improving models, is the major factor leading to dramatic ML model improvement. Within ML, model parameters are almost universally fit via gradient-based optimization with respect to well-defined objective functions, enabling seamless reproducibility, extensibility, and composability with models and fitting protocols made available via open-source platforms (e.g. GitHub, Hugging Face). These standardized and transparent fitting procedures enable rapid technological innovation [15, 16, 17, 18].
Comparatively, the accuracy of coarse-grained potentials has stagnated. Despite resounding successes in specific molecular models, force fields are often the product of “black magic” fitting procedures that are difficult to reproduce, typically involving significant hand design. Even relatively principled fitting procedures such as Iterative Boltzmann Inversion [19, 20], maximum likelihood methods [21], or Newton Inversion [22] are often performed piecemeal, with each interaction parameterized individually rather than performing a global optimization of the entire parameter set. As such, existing models cannot be easily adapted to account for new effects or incorporate new experimental data. A case in point is nucleic acid force field development, with (i) expanded parameterizations of DNA/RNA models appearing years after the original model release [23, 24, 25, 26, 27] and (ii) entirely new models developed to resolve existing shortcomings [20], attesting to the difficulty and laboriousness of refining models where parameterization is nontransparent and unsystematic. The ultimate result is a fragmented “zoo” of models that each require months or years of trial-and-error parameter fitting despite attempting to represent similar datasets.
Figure 1: Automatically fitting coarse-grained models to experimental data. A. For a given molecular system (e.g. from left to right, DNA, RNA, DNA-protein complexes), a loss function is defined that characterizes agreement between simulated macroscopic properties and experimental data. The gradient of this loss function is computed via automatic differentiation to update the parameters of the energy function. The depiction of the DNA and RNA models are adapted from Refs. [28] and [24]. B. An overview of differentiable Monte Carlo (or “DiffTRE”), a recently developed means of stochastic gradient estimation that operates at the level of unnormalized probability distributions [29, 30]. Reference states are sampled from the steady state distribution and the expected value of the target (time-independent) observable, $\langle O\rangle_{\theta}$, is reweighted via a set of differentiable weights. This provides a direct functional relationship between $\theta$ and $\langle O\rangle_{\theta}$ without differentiating through the simulation [29]; see Materials and Methods for details.
Here, we establish how coarse-grained, physics-based potentials can be fit to a broad class of experimental data with “frictionless reproducibility” (Figure 1A). We leverage differentiable programming, the backbone of machine learning algorithms whereby gradients are computed via automatic differentiation (AD). Traditional differentiable molecular dynamics (MD) is plagued by three core problems. First, computing the gradient through long unrolled simulations imposes significant memory overhead and numerical instabilities [31, 32]. Second, many fundamental algorithms in molecular simulation are not end-to-end differentiable [33, 34, 35]. Third, decades of research have been dedicated to optimizing MD algorithms to increase accessible timescales, but many of these optimizations are lost upon reimplementation in an automatic differentiation framework [36].
@@
We surmount these limitations using a novel method for stochastic gradient estimation [30, 29], making it possible to use gradient-descent to fit force field parameters to three broad classes of experimental data: structural, mechanical, and thermodynamic. Each class of data poses a unique set of challenges that are not present in toy models previously studied with differentiable MD. For example, calculation of mechanical properties, such as persistence length or stretch modulus, typically requires micro- and millisecond-scale simulations to sample many uncorrelated states, line-fitting procedures to infer summary statistics, and the application of external forces and torques [37, 24, 20, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Similarly, thermodynamic (e.g. melting temperature) calculations require particularly complex simulations, involving enhanced sampling, non-differentiable Monte Carlo algorithms, and temperature extrapolation [37, 24, 49, 50, 39, 38, 51, 52, 53, 54]. We devise frameworks for differentiating each such class of complex simulations. We also demonstrate how the calculated gradients can be adapted for sensitivity analyses and therefore the rapid derivation of physical insights.
Finally, we combine the methods developed for each individual class of data to fit parameters to single objective functions that jointly target structural, mechanical, and thermodynamic properties, allowing for the first time a systematic, simultaneous parameter optimization for different loss functions. We demonstrate how conflict-free gradients can be adapted from multi-task learning [55, 56, 57, 58, 59] to impose individual hard constraints despite expressing a global objective function.
In what follows, we demonstrate our approach by focusing first on an illustrative case study: nucleic acid force field development within oxDNA [60], a popular model of DNA that has enjoyed extensive experimental validation [37, 61, 62]. To reinforce the applicability of our approach to any coarse-grained model, we then detail sample extensions of our method to RNA and DNA-protein systems.
Figure 2: Parameter optimization for target structural properties via trajectory reweighting. A. A depiction of pitch, defined as the axial distance over which the helix makes one complete turn, and our loss function for fitting energy function parameters $\theta$ to achieve a target pitch. We define the loss as the root mean squared error between the simulated pitch $P^{\text{sim}}(\theta)$ and the target pitch $P^{\text{target}}$. B. The pitch as a function of iteration when using the gradients calculated via DiffTRE for gradient-based optimization. Iterations at which reference states were resampled are represented by scatter points. Varying only coaxial stacking parameters (orange line) fails because coaxial stacking – the stacking interaction between contiguous but distinct strands on a single side of a DNA duplex – was absent in these simulations. C. The ten most significant parameters ranked by the absolute value of their gradients. D. The gated DNA nanopore introduced in Ref. [63]. The structure is depicted both in PDB format and in the coarse-grained representation employed by oxDNA. E. The average RMSD of the simulated structure compared to the target structure when applying gradient-based optimization to maximize agreement between simulation and the ground-truth sequence/structure pair. This optimization uses the sequence-dependent variant of oxDNA2.
@@ 2.1 Differentiable Force Field Implementation
The first step towards gradient-based optimization of a molecular force field is implementing it in an AD library. We implement the oxDNA energy function in JAX, a state-of-the-art AD library for scientific computing [64], enabling the automatic calculation of arbitrary higher-order derivatives with respect to the control parameters. Unless otherwise stated, we focus our results on the latest sequence-averaged version of oxDNA, oxDNA2 [26]. The oxDNA2 model comprises 103 total parameters governing potential energy functions that capture excluded volume, base pairing, base stacking, and electrostatic interactions. These parameters were originally fit to structural, mechanical, and thermodynamic experimental data; we treat each in turn below. For further details of the model, see Materials and Methods. Using our implementation of the energy function, we can simulate systems of DNA using JAX-MD, an end-to-end differentiable molecular dynamics engine written in JAX [36]. We benchmark these simulations against the standalone oxDNA code [65, 66], involving both highly-optimized C++ and CUDA implementations, and find that simulation via JAX-MD is 2-3x slower (Figure S2) – a good compromise for the increased flexibility offered by the differentiable framework.
The JAX implementation offers two key advantages. First, since we write our energy function in an AD framework, forces are computed automatically from energy functions. This significantly reduces the barrier to modifying the functional form of the energy function and also permits the application of arbitrary (continuous and differentiable) bias potentials. Second, since the code can be compiled to a range of hardware accelerators, we can immediately run simulations on GPU and TPU without reimplementing the force field in a separate framework. Therefore, molecular simulation in an AD framework can be seen as complementary to highly optimized MD frameworks: beyond permitting gradient-based optimization, it enables flexibility and extensibility for force field development.
@@ 2.2 Structural Optimization
With a differentiable force field implementation in hand, we turn to optimizing parameters. Structural properties are generally the most amenable to optimization via differentiable MD since these properties converge even in relatively short simulations of small systems. As a first step, we fit a subset of oxDNA parameters to target pitch (Figure 2A) and propeller twist (Figure S4) values. We found that the two traditional means of stochastic gradient estimation for differentiable MD (Figure S1) were not sufficient for optimization in this case. Firstly, the pathwise estimator [67], whereby gradients are computed directly through an unrolled numerical simulation, imposes significant time and memory overhead and leads to exploding gradients [31]. The alternative, the score function estimator [67] – by which batches of simulations are performed and the probability of such simulations are increased or decreased according to the calculated observable – did not yield sufficient signal for optimization, likely due to unwieldy variance. See Figure S2 for benchmarking and Figure S3 for these optimization results with varying hyperparameters.
@@
We instead turn to a recently developed method of stochastic gradient estimation that directly operates at the level of unnormalized probability distributions (Figure 1B). This method was first introduced for differentiable Monte Carlo in the context of variational quantum Monte Carlo by Zhang et al. [30] and for machine-learned molecular simulation potentials [29], where it was termed “differentiable trajectory reweighting” (DiffTRE). The crux of the method is establishing a direct functional relationship between $\theta$, the control parameters of a distribution $p(x;\theta)$, and the expected value of a state-level observable $O$, $\mathbb{E}_{p(x;\theta)}[O(x;\theta)]$, that is independent of the sampling procedure. This is achieved via a perturbative calculation that expresses the expected value via a differentiable set of weights; see Materials and Methods for details. This estimator circumvents the memory and numerical issues associated with traditional differentiable MD by only differentiating the energies of reference states sampled from a set of simulations, rather than full trajectories.
@@
We find that for fitting structural properties to oxDNA, DiffTRE provides a numerically stable, scalable, and efficient method for optimization via gradient descent. Unlike the pathwise and score function estimators, this method yielded oxDNA parameters for specified pitch values within 50 iterations (Figure 2B), and the procedure was stable across various target pitch values. It is important to note that DiffTRE does not require that the sampling procedure itself is differentiable: reference states can be collected with simulations of any standalone MD code – in our case, the standalone oxDNA implementation [65].
Figure 3: Parameter optimization for target mechanical properties via trajectory reweighting. A. Our method for parameter optimization for complex target mechanical properties via DiffTRE and implicit derivatives. Individual ensembles are generated corresponding to the system’s response to a varying mechanical stimulus. In practice, multiple observables may be computed from a single ensemble (e.g. correlations at different distances in the passive calculation of persistence length). These values are then used to infer a summary statistic via an arbitrary line-fitting procedure. Individual observables are differentiated via DiffTRE [29] and line-fitting procedures are differentiated via implicit differentiation [68]. B. Parameter optimization for target values of persistence length ($L_{ps}$). I. A depiction of $L_{ps}$. II. Optimization of all oxDNA parameters to match various target $L_{ps}$ values. III. The top three most important parameters ranked via the absolute value of the gradient in the initial optimization step. C. Parameter optimization for target values of moduli characterizing response to external force and torque. I. A depiction of a duplex under external force and torque. II-III. Optimization of all oxDNA parameters to match the twist-stretch coupling, torsional modulus, and effective stretching modulus, all of which can be brought to experimental agreement (green bars) within error. Outliers in the twist-stretch coupling optimization at iterations 52 and 53 are omitted and replaced with a dashed line. See Figure S7 for corresponding sensitivity analyses. For B-C, iterations at which reference states were resampled are represented by scatter points.
@@
For a physics-based model, the gradient contains information about the most important physical interactions for a target property, accessible by ranking parameters based on the magnitude of their gradients (Figure 2C). In this work, we consider the gradient magnitudes in the first step of the optimization procedure, but more sophisticated analysis schemes are possible. For pitch, we find that the most significant parameter is the equilibrium distance of the backbone sites, as previously found [26]. More surprisingly, we find that cross-stacking interactions are quite important: this interaction is most often implicated for controlling duplex melting temperatures given its stabilizing role [69, 70, 71]. Sensitivity analysis for propeller twist (Figure S4) also reveals that the two most important parameters are involved in the cross-stacking interaction.
Finally, we can fit to macroscopic structural properties rather than those defined locally. As an example, we consider the gated DNA nanopore introduced in Ref. [63]. We minimize the root mean square deviation (RMSD) between the simulated structure and the ground-truth structure deposited in Nanobase [72], directly improving agreement between simulation and experiment (Figure 2D-E and Figure S5).